OpenStudy (anonymous):

How Do I Simplifiy 5/10 ? And If Your Gonna Be Rude, Dont Comment! Im Dislexic.

OpenStudy (anonymous):

It's 1/2.Just divide the 5 and 10 by 5. 5/5=1 & 10/5=2

OpenStudy (anonymous):

@TareKKowshiK, Thank You So Much For Helping Me! (:

OpenStudy (anonymous):

\[\frac{ 5 }{ 10 }=\frac{ 5/5 }{ 10/5 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Thanks @JAMG(:

OpenStudy (anonymous):

no problem always happy to help :D